OpenStudy (anonymous):

determine the best method to solve each system o equation.then solve the system. 1.4x=2y-10; x+2y=5 2.x=-2y; 4x+4y=-10 3.2s-3t=42; 3s+2t=24 4.4a-4b=-10; 2a+4b=-2 5.4x+10y=-6; -2-10y=2 6.2x=y-3; -x+y=0

OpenStudy (radar):

1. Go for substitution as you can let y=2x so you get in second equation x+4x=5 5x=5, x=1, y=2 2. Since x=-2y again substution. -8y+4y=-10, -4y=-10 y=5/2, x=-5

OpenStudy (radar):

1st equation times -3: -6s + 9t=-126 2nd equation times 2: 6s +4t = 48 add ------------- 13t =-78 t= -6 Now substitute -6 for t in first equation in 3. 2s+18=42, 2s=42-18=24 s=12

OpenStudy (radar):

I have done the first half, using both typical methods, you can continue with the last half.